OpenStudy (anonymous):

What is the equation in standard form of the link which passes through (6, 3) and has a slope of 2/3

OpenStudy (anonymous):

a. 2x + 3y = 3 b. 2x - 3y = 3 c. 3x + 2y = 3 d. 3x - 2y = 3

OpenStudy (akashdeepdeb):

Use the point-slope formula of a line to find this! Point-slope form = y-y1 = m(x-x1) You are given (x1,y1) = (6,3) m = 2/3 Put these values in the equation and get the equation with x and y as variables! :D Understood? :)

OpenStudy (anonymous):

Umm not really.. How exactly do I get the answer?